• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 08:20:09 on March 8,2007

    Entry number 195125

    keyword=shift summary

    00:00 MCC tuning beam in hall.
    00:40 10 uA in the hall.
    03:30 MCC calls. They can give us 70uA and let us take data while they figure out the beam conditions. But they want to run steady (e.g., postpone bullseye scan until later). Install target 4.

    03:36 Start run 1255, 75 uA.
    PS need to be adjusted.
    04:35 MCC calls to raise us to 80 uA.
    05:20 Raster horitzontal size is gone. MCC tries rebooting IOC but no improvement. MCC advises us to call Chris Cuevas or Bill Gunning in as techs-on-call for system. The pager number for Bill Gunning is not in service and other contact information is unavailable. Chris Cuevas does not answer his pager. Tried paging Chris twice. Call RC (A. Saha) who is now checking with MCC to figure out what we can do.

    05:40 MCC has 105 uA on target (raster off, empty target).